Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
May 2, 2008

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk
Announcements
  1. The chicken is dead at Bobbie's house. Cyrus the dog got him.
Review Minutes of 4/25/08
  1. Announcements: Two corrections in the Library Council recap: dealing with West reorg and departmental parking passes.
  2. Nets #3 - Changes made (past tense) to Domain Controller.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Orders have slowed - we're awaiting the budget announcement.
Liaisons
  1. Liaison training has been scheduled: MS certification for desktop support technician. It will take place the last two weeks of May.
Networking
  1. SX280 failures - half are out of warranty. Dell is working on having a tech replace all of the motherboards on the ones that are still under warranty and will send us motherboards to install on the out-of-warranty machines. This is all currently under negotiation.
  2. Monitor replacement - Evan spoke with Dell. They have agreed to replace 168 monitors. Several liaisons have been enlisted to help in the process. They should be arriving today in West. Bobbie will be supervising.
  3. McAfee - seems fine. There are one or two things that Bobbie is checking on.
  4. NT16 was full the other day. C: drive short on space - may need to make a virtual drive for printers - maybe next week.
  5. CJK - it's basically ready - Logan requested that the machine be brought to Systems for imaging.
  6. Two SX280s were out-of-warranty and broken. Cynthia replaced the motherboards. That repaired one, but the other was still not working. She made some adjustments and now they are both working. Need to re-swap the replacement given to Education.
Software Projects Status
  1. Wiki - Logan had a meeting on Monday with most of the interested participants. We decided to stick with Media Wiki. We will probably use it with the Circulation Manual. Technical Services has requested a Wiki. Logan will probably work on it next week. Secure wikis will need a certificate on the server. We will need one for the Web server as well. The Bridges affiliation will also require a certificate.
  2. Grover - testing, demo - this week: Orders & Networking.
  3. Tracking database - Tom is writing up a test document for changes.
  4. Greenstone - Mark is currently working in Chicago.
  5. Affiliation meeting last Friday - went well. Bridges had some questions asking us to justify why we need additional identifying information about the unaffiliated patrons that use our computers. Winston emailed Warren the info, but has not heard back yet.
  6. Winston got a request from FIU for the code that we use for Accounts Payable - the encrypted file from FCLA that we send to Bridges.
  7. Candidate assessment form is live.
DLC Greenstone Issues
  1. Scanner setup issues - may need to change the setup today. Logan is unsure of the cause. They keep having multiple time-outs and cannot save to the SAN. Could be the CNS switch - it may be bottlenecking at that point. May save scans to a central location and then push to the SAN overnight. DLC wants to upgrade one or two ports. Unsure how to proceed with that, given the current budget. Bobbie to call Mary at CNS to find out the cost for a 1G switch.
Web Projects
  1. Last night Tom put up the survey for extended hours. He basically just changed the dates from last semester's survey.
  2. Library videos - Quicktime & Windows media format - 3.5G - Tom needs to know where to put it on the Web server. Jason Fleming and Laurie Taylor need to be able to access it. We will put them on the DLCAV server. They could also go up on YouTube.
Students and Departments
  1. Ad expired. New one had to be created.
  2. Lila and Rob are working through Thursday. Luiza's last day is today.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page